OpenStudy (anonymous):

algebra help please will fan and give best answer to fastest and most accurate answer :-) Stephanie is organizing the seating arrangements for a concert in her local park. The first row's seats are modeled by the function f(x) = 15. To give everyone a good view, each row must have four more seats than the row before it. The number of additional seats a row has is modeled by the function s(x) = 4(x - 1). Explain to Stephanie how to create an equation to predict the number of seats in any row. Describe how to determine the number of seats in the fortieth row. Use complete sentences.

OpenStudy (anonymous):

@ganeshie8 @amistre64 @Shay17 @jacob4900

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@jim_thompson5910 please anyone help me

OpenStudy (anonymous):

hold on let me solve it.

OpenStudy (anonymous):

thank you @jacob4900

OpenStudy (anonymous):

no problem. i just hope that i can solve it because i haven't studied this in a while

hero (hero):

Row 1 = 15 Row 2 = 15 + 4 = 19 Row 3 = 15 + 4 + 4 = 23 Row 4 = 15 + 4 + 4 + 4 = 27 Row 5 = 15 + 4 + 4 + 4 + 4 = 31 Row n = `15 + 4(n - 1)`

hero (hero):

Or better yet: Row 1 = 15 Row 2 = 15 + 4(1) = 19 Row 3 = 15 + 4(2) = 23 Row 4 = 15 + 4(3) = 27 Row 5 = 15 + 4(4) = 31 Row n = 15 + 4(n - 1)

hero (hero):

Now find the number of seats in Row 40

OpenStudy (anonymous):

11 +4x where x is row number For fortieth row...171 seats

OpenStudy (anonymous):

ok @hero so n=15+4(40-1) n=15+4(39) n=15+156 n=171 like @tanmaykhandait said

OpenStudy (anonymous):

did i do that right?

hero (hero):

@tanmaykhandait just simplified what I already wrote: 15 + 4(n - 1) = 15 + 4n - 4 = 15 - 4 + 4n = 11 + 4n

OpenStudy (anonymous):

okay i understand so did i do it correctly?

hero (hero):

I used n instead of x because n is usually used for these kinds of problems

OpenStudy (anonymous):

i understand and im awrding best answer but can you help me with a few more?

hero (hero):

But really it simplifies to just f(x) + s(x)